# Formal conjugacy and asymptotic differential algebra

### v3

@@ 3.1 Transseries
Let $\mathbb{T}$ denote the ordered field of logarithmic-exponential transseries [16, 12] together with its standard derivation $\partial:f\mapsto f^{\prime}$ and let $\mathfrak{M}$ denote the set of transmonomials in $\mathbb{T}$, i.e. $\mathfrak{M}$ is a specific section of the valuation $v:\mathbb{T}\longrightarrow\Gamma\cup\{\infty\}$. We identify each value $\gamma\in v(\mathbb{T}^{\times})$ with the corresponding transmonomial $\mathfrak{m}\in\mathfrak{M}$ with $v(\mathfrak{m})=\gamma$. Recall that $\mathbb{T}$ is a direct limit of Hahn fields $\mathbb{T}_{m,n},m,n\in\mathbb{N}$ where $\mathbb{T}_{m,n}=\mathbb{R}\left(\!\!(\mathfrak{M}_{m,n})\!\!\right)$ is the field of transseries with exponential depth $\leqslant\>m$ and logarithmic depth $\leqslant\>n$ (see [16, 18]). The derivation $\partial$ is regular [12, Section 3]. The identity series is denoted $x$, and we have $x^{\prime}=1$. There is a bijective morphism $\log:(\mathbb{T}^{>},\cdot,1,<)\longrightarrow(\mathbb{T},+,0,<)$ where $\mathbb{T}^{>}=\{f\in\mathbb{T}:f>0\}$. For each $n\in\mathbb{N}$, we denote by $\log_{n}x$ the $n$-th iterate of $\log$ applied at $x$. Set $\mathbb{T}^{>\mathbb{R}}:=\{f\in\mathbb{T}:f>\mathbb{R}\}$. We recall that $\mathbb{T}$ is equipped with a formal composition law $\mathord{\circ}:\mathbb{T}\times\mathbb{T}^{>\mathbb{R}}\longrightarrow\mathbb{T}$ (see [12, Section 6]). We write $\check{\circ}$ for the inverse law on $\mathbb{T}^{>\mathbb{R}}$, given by $f\check{\circ}g:=g\circ f$ for all $f,g\in\mathbb{T}^{>\mathbb{R}}$.
Let $\mathbb{S}$ denote the subset of $\mathbb{T}$ of flat transseries, i.e. series $f\in\mathbb{T}$ with $\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}$ for all $\mathfrak{m}\in\operatorname{supp}f$. So $\mathbb{S}=\{f\in\mathbb{T}:\operatorname{supp}f\subseteq\mathfrak{M}_{\preceq\!\!\!\preceq x}\}$ where $\mathfrak{M}_{\preceq\!\!\!\preceq x}=\{\mathfrak{m}\in\mathfrak{M}:\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}\}$. Note that $\mathfrak{M}_{\preceq\!\!\!\preceq}$ is a subgroup of $\mathfrak{M}$, so $\mathbb{S}$ is a subfield of $\mathbb{T}$. We write $\mathbb{S}^{>\mathbb{R}}=\mathbb{S}\cap\mathbb{T}^{>\mathbb{R}}$. We have $\partial(\mathfrak{M}_{\preceq\!\!\!\preceq x})\subseteq\mathbb{S}$ as a consequence of our results on near supports of derivations over transseries [5, Theorem 6.7], so $\mathbb{S}$ is a differential subfield of $\mathbb{T}$. Likewise, we have $\mathfrak{M}_{\preceq\!\!\!\preceq x}\circ f\subseteq\mathbb{S}$ for all $f\in\mathbb{S}^{>\mathbb{R}}$ as a consequence of our results on relative near-supports [5, Proposition 7.31 and Theorem 7.1]. We now focus on the H-field $(\mathbb{S},\preccurlyeq,\partial)$, and the group $(\operatorname{Cont}(\partial),\ast,0)$.
@@
The direct limit $\mathbb{S}$ has regular asymptotic integration.
@@
Let $(m,n)\in\mathbb{N}^{2}$ and $f\in\mathbb{S}_{m,n}$. The pseudo-gap of $\mathbb{S}_{m,n}$ is $\gamma_{n}:=v((\log_{n+1}x)^{\prime})$ where $\log_{n+1}x\in\mathbb{S}_{m,n+1}\setminus\mathbb{S}_{m,n}$. Thus if $v(f)\neq\gamma_{n}$, then $f$ has an asymptotic integral in $\mathbb{S}_{m,n}$. If $v(f)=\gamma_{n}$, then $c\log_{n+1}x$ is an asymptotic integral of $f$ in $\mathbb{S}_{m,n+1}$ with derivative in $\mathbb{S}_{m,n}$, where $c\in C$ is the leading coefficient of $f$. We also deduce that there is no pseudo-gap in $\mathbb{S}$, so the conditions of Definition 1.2 hold for $d_{1}=(m,n+1)$. ∎
@@
Right compositions satisfy chain rules [12, Proposition 6.3]. In fact, these notions coincide:
@@
Let $\sigma\in 1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ satisfy a chain rule. Then $\sigma$ is a right composition.
@@
Consider the contracting map $\phi:=\sigma-\operatorname{Id}_{\mathbb{S}}$. The chain rule condition Equation 5 is preserved under composition with regular automorphisms satisfying the chain rule. So conjugating $\sigma$ by the right composition with $x+x^{-1}$, we may assume that $\phi(\mathbb{S})\subseteq\mathbb{S}^{\prec}$.
Note that $\log(\mathfrak{M}_{\preceq\!\!\!\preceq x})\subseteq\mathfrak{M}_{\preceq\!\!\!\preceq x}$ since $1\prec\operatorname{supp}\log\mathfrak{m}\preccurlyeq\mathfrak{m}$ for all $\mathfrak{m}\in\mathfrak{M}$ and $\mathbb{T}$ has H-type. Thus $\log(\mathbb{S}\cap\mathbb{T}^{>})\subseteq\mathbb{S}$. Consider an $s\in\mathbb{S}^{>\mathbb{R}}$. We have
@@
$=$ $0$
So $c:=\log(\sigma(s))-\sigma(\log s)\in\operatorname{Ker}(\partial)=C$. We have
$\log s+\phi(\log s)=\sigma(\log s)=\log(\sigma(s))-c=\log(s+\phi(s))-c=\log s+\delta-c$
where $\varepsilon:=\log(s+\phi(s))-\log s\sim\frac{\phi(s)}{s}\prec 1$ and $\phi(\log s)\prec 1$. So $c\prec 1$. But $c\in C$, so $c=0$, i.e. $\log(\sigma(s))=\sigma(\log s)$. Since this holds for all $s\in\mathbb{S}^{>\mathbb{R}}$, we deduce that $\sigma$ must be a right composition, hence the result. ∎
@@
The derivation $\partial$ is contracting on $\mathbb{S}$ and $\exp(\partial)$ coincides with the right composition with $x+1$.
@@
Now $1/V^{\prime}x\prec 1$ so $(V^{\operatorname{inv}})^{{\dagger}}\prec 1$ by [24, Proposition 5.10].
We claim that $V^{{\dagger}}\preccurlyeq x^{-1}$. Indeed, we have $f-x>x^{-n}$ for a certain $n>1$. Note that $x^{n}\circ(x+x^{-n})>x^{n}+1$. The ordered group $(\mathbb{T}^{>\mathbb{R}},\circ,x,<)$ is a growth order group with Archimedean centralisers as a consequence of [8, Theorem 4.7]. So by the axiom GOG2 of [8, Section 2.1], for all $\varphi,\psi\in\mathbb{T}^{>\mathbb{R}}$ such that $\psi$ lies above all iterates of $x+x^{-n}$ and $\varphi>\psi\circ\psi$, we have $\varphi\circ(x+x^{-n})\circ\varphi^{\operatorname{inv}}>\psi\circ(x+x^{-n})\circ\psi^{\operatorname{inv}}$. Here we apply this to $\psi=x^{n}$, and see that $V\leqslant x^{n}$ for a certain $n\in\mathbb{N}$. Thus $V^{{\dagger}}\preccurlyeq x^{-1}$. Now writing $V=V_{0}+V_{1}$ where $V_{0}\in\mathbb{S}$ and $V_{1}^{{\dagger}}\succ x^{-1}$, we have $V_{0}\circ f+V_{1}\circ f=V_{0}+1+V_{1}+1$ where $V_{0}\circ f-V_{0}-1\in\mathbb{S}$, so we must have $V_{1}\circ f-V_{1}=0$, and thus we may assume that $V=V_{0}\in\mathbb{S}$. By Lemma 3.4, we have
@@
We have $\frac{1}{V^{\prime}}\prec x$, so the regular derivation $\mathrm{d}:=\frac{1}{V^{\prime}}\partial$ on $\mathbb{S}$ is contracting by Lemma 3.2. An easy induction using the chain rule [12, Proposition 6.3] shows that $(V^{\operatorname{inv}})^{(i)}\circ V$ is the value of the $i$-th iterate of $\mathrm{d}$ at $x$, for all $i>0$. So $\exp(\mathrm{d})(x)=f$. Recall that $\exp(\mathrm{d})$ satisfies a chain rule with respect to $\partial$, thus by Lemma 3.3 it is the right composition with $\exp(\mathrm{d})(x)=f$. Now $\exp:\operatorname{Der}^{\mathcal{S}}(\mathbb{S})\longrightarrow 1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ is a group morphism, and $\sigma\mapsto\sigma(x)$ is a group morphism $\left(1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S}),\circ,\operatorname{Id}_{\mathbb{S}}\right)\longrightarrow(\mathcal{P},\check{\circ},x)$, so the result follows. ∎
@@
Note that $\exp(f\partial)(x)-x\sim f$ for all $f\in\operatorname{Cont}(\partial)$. Recall that $v(x^{-1})=\min\Psi$. In view of Lemma 3.1, we may apply Theorems 3.1 and 2.1, and obtain the conditions $\varepsilon\sim\delta$ and $v\left(\frac{\varepsilon-\delta}{\delta^{2}}\right)>v\left(\left(\frac{x}{\delta}\right)^{\prime}\right)$. Since $1-\delta x^{-1}\preccurlyeq 1$, the inequality $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$ entails that $\varepsilon\sim\delta$. We conclude with Lemma 3.5. ∎
@@
    $\partial(\mathbb{U})\subseteq\mathbb{U}$ and $\partial$ is $\mathcal{U}$-regular, and
@@
For instance, the direct system $(\mathfrak{M}_{0,n})_{n\in\mathbb{N}}$ corresponding to the field $\mathbb{T}_{\log}$ of logarithmic transseries [21], is a subsystem of $\mathcal{S}$.
@@
The same arguments as in Lemma 3.1 show that $\mathbb{T}_{\log}$ has regular asymptotic integration. Now the set $v(\mathbb{T}_{0,n})$ lies above the poincaré factor $v((\log_{n+1})^{\prime})$ of $\mathbb{T}_{0,n+1}$. So as in Theorem 3.2, we may apply Proposition 3.1 and Proposition 2.1, and conclude. ∎
@@
Consider the subsystem $\left(\frac{1}{n!}\mathbb{Z}\right)_{m,n\in\mathbb{N}}$ of $\mathcal{S}$. This corresponds to formal Puiseux series over $\mathbb{R}$. Here there is resonance, so we only obtain the resonance-free sufficient but non-necessary conditions as in Theorem 3.3.

### v4

@@ 3.1 Transseries
Let $\mathbb{T}$ denote the ordered field of logarithmic-exponential transseries [16, 12] together with its standard derivation $\partial:f\mapsto f^{\prime}$ and let $\mathfrak{M}$ denote the set of transmonomials in $\mathbb{T}$, i.e. $\mathfrak{M}$ is a specific section of the valuation $v:\mathbb{T}\longrightarrow\Gamma\cup\{\infty\}$. We identify each value $\gamma\in v(\mathbb{T}^{\times})$ with the corresponding transmonomial $\mathfrak{m}\in\mathfrak{M}$ with $v(\mathfrak{m})=\gamma$. Recall that $\mathbb{T}$ is a direct limit of Hahn fields $\mathbb{T}_{m,n},m,n\in\mathbb{N}$ where $\mathbb{T}_{m,n}=\mathbb{R}\left(\!(\mathfrak{M}_{m,n})\!\right)$ is the field of transseries with exponential depth $\leqslant m$ and logarithmic depth $\leqslant n$ (see [16, 18]). The directed system $(\mathfrak{M}_{m,n})_{m,n\in\mathbb{N}}$ has inclusions $\mathfrak{M}_{m,n}\longrightarrow\mathfrak{M}_{m^{\prime},n^{\prime}}$ whenever $n^{\prime}\geqslant n$ and $m^{\prime}\geqslant m+(n^{\prime}-n)$ (see [12, (2.7)]). The derivation $\partial$ is regular [12, Section 3]. The identity series is denoted $x$, and we have $x^{\prime}=1$, so $x^{{\dagger}}=x^{-1}$. Note that contrary to [31, 32, 30, 29], we take $x$ to be positive infinite as opposed to positive infinitesimal. There is a bijective morphism $\log:(\mathbb{T}^{>},\cdot,1,<)\longrightarrow(\mathbb{T},+,0,<)$ where $\mathbb{T}^{>}=\{f\in\mathbb{T}:f>0\}$. Given $n\in\mathbb{N}$, we denote by $\log_{n}x$ the $n$-th iterate of $\log$ applied at $x$. We write $\mathbb{T}_{\log}:=\bigcup\limits_{n\in\mathbb{N}}\mathbb{T}_{0,n}$ for the field of logarithmic transseries [21].
[…]
Let $\mathbb{S}$ denote the subset of $\mathbb{T}$ of flat transseries, i.e. series $f\in\mathbb{T}$ with $\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}$ for all $\mathfrak{m}\in\operatorname{supp}f$. So $\mathbb{S}=\{f\in\mathbb{T}:\operatorname{supp}f\subseteq\mathfrak{M}_{\preceq\!\!\preceq x}\}$ where $\mathfrak{M}_{\preceq\!\!\preceq x}=\{\mathfrak{m}\in\mathfrak{M}:\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}\}$. Note that $\mathfrak{M}_{\preceq\!\!\preceq}$ is a subgroup of $\mathfrak{M}$, so $\mathbb{S}$ is a subfield of $\mathbb{T}$. We write $\mathbb{S}^{>\mathbb{R}}:=\mathbb{S}\cap\mathbb{T}^{>\mathbb{R}}$ and $\mathbb{S}^{>}:=\mathbb{S}\cap\mathbb{T}^{>}$.
Example 3.1
The transmonomials $(\log_{2}x)^{-5}$, $x^{2}$ and $\operatorname{e}^{\sqrt{\log x}}$ are flat, whereas $\operatorname{e}^{(\log x)^{2}}$ and $\operatorname{e}^{x}$ are not.
@@
$\mathbb{S}$ is a differential subfield of $\mathbb{T}$
@@
It suffices to show that $\partial(\mathfrak{M}_{\preceq\!\!\preceq x})\subseteq\mathbb{S}$. We prove by induction on $m\in\mathbb{N}$ with $\mathfrak{m}\in\bigcup\limits_{n\in\mathbb{N}}\mathbb{T}_{m,n}$ that $\mathfrak{m}\in\mathfrak{M}_{\preceq\!\!\preceq x}\Rightarrow\mathfrak{m}^{\prime}\in\mathbb{S}$. This is clear if $m=0$ since $\mathbb{T}_{\log}^{\prime}\subseteq\mathbb{T}_{\log}$. If $\mathfrak{m}\in\mathbb{T}_{m+1,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x}$ and the result holds at $m$, then we have $\operatorname{supp}\mathfrak{m}^{\prime}\subseteq\mathfrak{m}\operatorname{supp}\log(\mathfrak{m})^{\prime}$ where $\log(\mathfrak{m})\in\mathbb{T}_{m,n+1}$. We conclude by induction.∎
Lemma 3.2
$\mathbb{S}$ is closed under composition.
Proof.
Let $f\in\mathbb{S}$ with $f>\mathbb{R}$. It suffices to show that $\mathfrak{M}_{\preceq\!\!\preceq x}\circ f\subseteq\mathbb{S}$. We prove by induction on $m\in\mathbb{N}$ with $\mathfrak{m}\in\bigcup\limits_{n\in\mathbb{N}}\mathbb{T}_{m,n}$ that $\mathfrak{m}\in\mathfrak{M}_{\preceq\!\!\preceq x}\Rightarrow\mathfrak{m}\circ f\in\mathbb{S}$. The group $\mathfrak{M}_{\preceq\!\!\preceq x}$ is closed under taking real powers, whence by [12, 6.8] so is $\mathbb{S}$. Moreover $\mathbb{S}$ is closed under $\log$. So the result holds if $m=0$. Suppose that $\mathfrak{m}\in\mathbb{T}_{m+1,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x}$ and that the result holds at $m$. Considering $\mathfrak{m}^{-1}$ if necessary, we may assume that $\mathfrak{m}$ is infinite. We have
$\mathfrak{m}\circ f=\exp(\log(\mathfrak{m})\circ f).$
By construction of $\exp$ on $\mathbb{T}$ (see [12, 1.5–2.7]), each element $\mathfrak{n}$ in $\operatorname{supp}\mathfrak{m}\circ f$ has the form $\mathfrak{n}=\mathfrak{u}\mathfrak{n}_{1}\cdots\mathfrak{n}_{k}$ where $\mathfrak{u}$ is the largest monomial of $\exp(\log(\mathfrak{m})\circ f)$, $k\in\mathbb{N}$, and $\mathfrak{n}_{1},\dots,\mathfrak{n}_{k}$ are in $\operatorname{supp}\log(\mathfrak{m})\circ f$. As $\log\mathfrak{m}\in\mathbb{T}_{m,n+1}$, the induction hypothesis gives $\mathfrak{n}_{1}\cdots\mathfrak{n}_{k}\in\mathbb{S}$. Lastly, note that $\mathfrak{u}^{{\dagger}}\asymp(\mathfrak{m}\circ f)^{{\dagger}}$. The chain rule [12, Proposition 6.3] gives $(\mathfrak{m}\circ f)^{{\dagger}}=f^{\prime}\mathfrak{m}^{{\dagger}}\circ f$. But $\mathfrak{m}^{{\dagger}}\preccurlyeq x^{-1}$ and the composotion by $f$ on the right is valuation preserving, so $(\mathfrak{m}\circ f)^{{\dagger}}\preccurlyeq f^{{\dagger}}\preccurlyeq x^{-1}$. So $\mathfrak{u}\in\mathbb{S}$, whence $\mathfrak{m}\circ f\in\mathbb{S}$. We conclude by induction. ∎
We have a directed system $\mathcal{S}:=(\mathfrak{M}_{m,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x})_{m,n\in\mathbb{N}}$ and we write $\mathbb{S}_{m,n}=\mathbb{R}(\!(\mathfrak{M}_{m,n}\cap\mathfrak{M}_{\preceq\!\!\preceq x})\!)=\mathbb{S}\cap\mathbb{T}_{m,n}$ for all $m,n\in\mathbb{N}$. We now focus on the directed system $\mathcal{S}$, H-field $(\mathbb{S},\preccurlyeq,\partial)$, and the group $(\operatorname{Cont}(\partial),\ast,0)$.
Lemma 3.3
The direct limits $\mathbb{S}$, $\mathbb{T}$ and $\mathbb{T}_{\log}$ have regular asymptotic integration.
Proof.
We prove this for $\mathbb{S}$, the proofs in the other cases being identical. Let $(m,n)\in\mathbb{N}^{2}$ and $f\in\mathbb{S}_{m,n}$. The pseudo-gap of $\mathbb{S}_{m,n}$ is $\gamma_{n}:=v((\log_{n+1}x)^{\prime})$ where $\log_{n+1}x\in\mathbb{S}_{m,n+1}\setminus\mathbb{S}_{m,n}$. Thus if $v(f)\neq\gamma_{n}$, then $f$ has an asymptotic integral in $\mathbb{S}_{m,n}$. If $v(f)=\gamma_{n}$, then $c\log_{n+1}x$ is an asymptotic integral of $f$ in $\mathbb{S}_{m,n+1}$ with derivative in $\mathbb{S}_{m,n}$, where $c\in C$ is the leading coefficient of $f$. We also deduce that there is no pseudo-gap in $\mathbb{S}$, so the conditions of Definition 1.2 hold for $(d_{0},d)=((m,n),(m,n+1))$. ∎
@@
Right compositions satisfy chain rules [12, Proposition 6.3]. We have a converse relation:
@@
Let $\sigma\in 1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ satisfy a chain rule. If $\sigma(\log s)-\log s\prec 1$ for all $s\in\mathbb{S}^{>}$, then $\sigma$ is the right composition with $\sigma(x)$.
@@
Consider the contracting map $\phi:=\sigma-\operatorname{Id}_{\mathbb{S}}$. Let $\mathfrak{m}\in\mathfrak{M}$ with $\mathfrak{m}\succ 1$ and let $\mathfrak{n}\in\operatorname{supp}\log\mathfrak{m}$. We have $1\prec\mathfrak{n}\prec\mathfrak{m}$ by construction of $\log$ on $\mathbb{T}$. It follows since $(\mathbb{T},\preccurlyeq,\partial)$ is an H-asymptotic field that $\mathfrak{n}^{{\dagger}}\asymp(\mathfrak{n}^{-1})^{{\dagger}}\preccurlyeq(\mathfrak{m}^{-1})^{{\dagger}}\asymp\mathfrak{m}^{{\dagger}}$. This entails that $\log(\mathfrak{M}_{\preceq\!\!\preceq x})\subseteq\mathfrak{M}_{\preceq\!\!\preceq x}$, so $\log\mathbb{S}^{>}\subseteq\mathbb{S}$. Consider an $s\in\mathbb{S}^{>\mathbb{R}}$. We have
@@
$=$ $0.$
So $c:=\log(\sigma(s))-\sigma(\log s)\in\operatorname{Ker}(\partial)=C$. Our hypothesis on $\sigma$ gives $c\prec 1$, whence $c=0$, i.e. $\log(\sigma(s))=\sigma(\log s)$. An easy induction on the directed system $(\mathbb{S}_{m,n})_{m,n\in\mathbb{N}}$ shows that the right composition with $\sigma(x)$ is the only strongly linear automorphism of $\mathbb{S}$ which commutes with $\log$ and sends $x$ to $\sigma(x)$. So it must coincide with $\sigma$. ∎
@@
If $g\in\operatorname{Cont}(\partial)$, then $\exp(g\partial)$ is a right composition. Moreover, the derivation $\partial$ is contracting on $\mathbb{S}$ and $\exp(\partial)$ coincides with the right composition with $x+1$.
@@
Note that $\sigma:=\exp(g\partial)$ commutes with $g\partial$. Thus
$\partial\circ\sigma=\frac{1}{g}(g\partial)\circ\sigma=\frac{1}{g}\sigma\circ(g\partial)=\frac{\sigma(g)}{g}\sigma\circ\partial.$
Therefore $\sigma$ satisfies a chain rule.
Let $s\in\mathbb{S}^{>}$. We have $\sigma(\log s)-\log s\sim g(\log s)^{\prime}$. Since $g\in\operatorname{Cont}(\partial)$, we have $g(\log s)^{\prime}\prec(\log s)^{\prime}=s^{{\dagger}}$. But $s\in\mathbb{S}^{\times}$, so $s^{{\dagger}}\preccurlyeq x^{-1}$, so $\sigma(\log s)-\log s\prec 1$. Applying the previous lemma, we see that $\sigma$ is a right composition.
@@
Now $1/(V^{\prime}x)\prec 1$ so $(V^{\operatorname{inv}})^{{\dagger}}\prec 1$ by [24, Proposition 5.10].
We claim that $V^{{\dagger}}\preccurlyeq x^{-1}$. Indeed, we have $f-x>x^{-n}$ for a certain $n>1$. Note that $x^{n+1}\circ(x+x^{-n})>x^{n+1}+1$. The ordered group $(\mathbb{T}^{>\mathbb{R}},\circ,x,<)$ is a growth order group with Archimedean centralisers as a consequence of [8, Theorem 4.7]. So by the axiom GOG2 of [8, Section 2.1], for all $\varphi,\psi\in\mathbb{T}^{>\mathbb{R}}$ such that $\psi$ lies above all iterates of $x+x^{-n}$ and that $\varphi>\psi\circ\psi$, we have $\varphi\circ(x+x^{-n})\circ\varphi^{\operatorname{inv}}>\psi\circ(x+x^{-n})\circ\psi^{\operatorname{inv}}$. Here we apply this to $\psi=x^{n+1}$, and see that we must have $V\leqslant x^{2(n+1)}$. Thus $V^{{\dagger}}\preccurlyeq x^{-1}$. Now writing $V=V_{0}+V_{1}$ where $V_{0}\in\mathbb{S}$ and $V_{1}^{{\dagger}}\succ x^{-1}$, we have $V_{0}\circ f+V_{1}\circ f=V_{0}+1+V_{1}+1$ where $V_{0}\circ f-V_{0}-1\in\mathbb{S}$, so we must have $V_{1}\circ f-V_{1}=0$, and thus we may assume that $V=V_{0}\in\mathbb{S}$. By Lemma 3.6, we have
@@
We have $\frac{1}{V^{\prime}}\prec x$, so the regular derivation $\mathrm{d}:=\frac{1}{V^{\prime}}\partial$ on $\mathbb{S}$ is contracting by Lemma 3.4. An easy induction using the chain rule [12, Proposition 6.3] shows that $(V^{\operatorname{inv}})^{(i)}\circ V$ is the value of the $i$-th iterate of $\mathrm{d}$ at $x$, for all $i>0$. So $\exp(\mathrm{d})(x)=f$. We see with Lemma 3.6 that $\exp(\mathrm{d})$ is the right composition with $f$. This shows that $\mathcal{E}$ is surjective. Now $\exp:\operatorname{Der}^{\mathcal{S}}(\mathbb{S})\longrightarrow 1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S})$ is a group morphism, and $\sigma\mapsto\sigma(x)$ is a group morphism $\left(1\text{-}\operatorname{Aut}^{\mathcal{S}}(\mathbb{S}),\circ,\operatorname{Id}_{\mathbb{S}}\right)\longrightarrow(\mathcal{P},\check{\circ},x)$, so the result follows. ∎
@@
Note that $\exp(f\partial)(x)-x\sim f$ for all $f\in\operatorname{Cont}(\partial)$. Recall that $v(x^{-1})=\min\Psi$. By Lemma 3.3, the direct limit $\mathbb{S}$ has regular asymptotic integration, so Theorem 2.1 applies. Now by Theorem 3.1 we can translate the conditions of conjugacy in Theorem 2.1, as the conjunction $\varepsilon\sim\delta$ and $v\left(\frac{\varepsilon-\delta}{\delta^{2}}\right)>v\left(\left(\frac{x}{\delta}\right)^{\prime}\right)$. Since $1-\delta x^{-1}\preccurlyeq 1$, the inequality $\varepsilon-\delta\prec\delta(1-x\delta^{{\dagger}})$ entails that $\varepsilon\sim\delta$. We conclude with Lemma 3.7. ∎
@@
    $\partial$ is $\mathcal{U}$-regular, and
@@
For instance, the direct system $(\mathfrak{M}_{0,n})_{n\in\mathbb{N}}$ corresponding to the field $\mathbb{T}_{\log}$ of logarithmic transseries, is a subsystem of $\mathcal{S}$.
@@
Recall by Lemma 3.3 that $\mathbb{T}_{\log}$ has regular asymptotic integration. The pseudo-gap $v((\log_{n+1}x)^{\prime})$ of $\mathbb{T}_{0,n+1}$ is strictly below the set $v(\mathbb{T}_{0,n})$. So we may apply Proposition 2.1 in $\operatorname{Cont}(\partial\restriction\mathbb{T}_{0,n+1})$. Then, as in the proof of Theorem 3.2, we translate the inequalities using the isomorphism from Proposition 3.1. ∎
@@
Consider the subsystem $\left(\frac{1}{n!}\mathbb{Z}\right)_{m,n\in\mathbb{N}}$ of $\mathcal{S}$. This corresponds to formal Puiseux series over $\mathbb{R}$. Here we have a pseudo-gap $v(x^{-1})$, so we only obtain the sufficient but non-necessary conditions as in Theorem 3.3.
Remark 3.2
Our results up to and excluding Theorem 3.1 apply when replacing $\mathbb{T}$ with the larger, class-sized field of $\omega$-series [5, Section 4.5]. This is an totally ordered direct limit of (set-sized) fiels of generalised series indexed by all ordinals. Although it has a pseudo-gap, asymptotic integration is regular with respect to this directed system.